Problem 8. Exhibit an isometry between the spaces C(0,1) and C(1,2)
Answer #1

We need to define a bijective function $\zeta:\mathcal{C}_{[1,2]}\to\mathcal{C}_{[0,1]}$ such that $d_{\infty}(f,g)=d_{\infty}(\zeta(f),\zeta(g))$ .

Define $\zeta:\mathcal{C}_{[1,2]}\to\mathcal{C}_{[0,1]}$ by $\zeta(f)=f(x+1)$ .

Then $\zeta^{-1}:\mathcal{C}_{[0,1]}\to\mathcal{C}_{[1,2]}$ is given by $\zeta^{-1}(f)=f(x-1)$ .

Now,

d_{\infty}(\zeta(f),\zeta(g))=sup_{x\in[0,1]}|\zeta(f)-\zeta(g)|\\=sup_{x\in[0,1]}|f(x+1)-g(x+1)|\\=sup_{x\in[1,2]}|f(x)-g(x)|\\=d_{\infty}(f,g)

Hence, \mathcal{C}_{[1,2]} and \mathcal{C}_{[0,1]} are isometric.